In class we showed the following: n ∑(k=1) k = n(n+1)/2 and n∑(k=1) k^2 = n(n+ 1)(2n+ 1)/6
Using the fact that (k+ 1)^4-k^4= 4k^3+ 6k^2+ 4k+ 1 and summing up as k= 1,2,3,……, n together with the above two equalities, deduce that n∑(k=1) k^3 = (n(n+1)/2)^2
1
Expert's answer
2020-10-13T19:31:22-0400

Let

"S_1 =\sum\limits_{k=1}^n1 = 1+1+1+...+1 = 1\cdot n = n""S_2 =\sum\limits_{k=1}^nk=\frac{n(n+1)}{2}""S_3 = \sum\limits_{k=1}^nk^2=\frac{n(n+1)(2n+1)}{6}""S_4 = \sum\limits_{k=1}^nk^3"

Use the fact


"(k+1)^4 -k^4 =4k^3 +6k^2 +4k+1"

Put "k=1,2,3..., n" to this formula:

"2^4 -1^4 =4\cdot 1^3+6\cdot 1^2 +4\cdot 1+1"

"3^4 -2^4 =4\cdot 2^3+6\cdot 2^2 +4\cdot 2+1"

"4^4 -3^4 =4\cdot 3^3+6\cdot 3^2 +4\cdot 3+1"

...

"(n+1)^4 -n^4 =4\cdot n^3+6\cdot n^2 +4\cdot n+1"

Sum left sides of these equalities:

"\cancel{2^4} -1^4 +\cancel{3^4} -\cancel{2^4} +\cancel{4^4} -\cancel{3^4} +...+(n+1)^4 -\cancel{n^4}=(n+1)^4-1"

Sum right sides of these equalities:

"4(1^3+2^3+...+n^3)+6(1^2+2^2+...+n^2)+4(1+2+...+n)+(1+1+...+1)="

"=4S_4+6S_3 +4S_2 +S_1"

Hence

"4S_4+6S_3 +4S_2 +S_1 =(n+1)^4-1"

"4S_4=(n+1)^4-1-6S_3-4S_2-S_1"

"4S_4 =(n+1)^4 -1-\frac{6n(n+1)(2n+1)}{6}-\frac{4n(n+1)}{2}-n"

"4S_4 =(n+1)^4-1-n(n+1)(2n+1)-2n(n+1)-n"

"4S_4 =(n+1)^4-n(n+1)(2n+1)-2n(n+1)-(n+1)"

"4S_4 =(n+1)((n+1)^3 -n(2n+1)-2n-1)"

"4S_4 =(n+1)(n^3+3n^2+\cancel{3n}+\cancel{1}-2n^2-\cancel{n}-\cancel{2n}-\cancel{1})"

"4S_4 =(n+1)(n^3+n^2)"

"4S_4 =(n+1)\cdot n^2(n+1)"

"4S_4 =n^2(n+1)^2"

"S_4 =\frac{n^2(n+1)^2}{4}=(\frac{n(n+1)}{2})^2"

"\sum\limits_{k=1}^nk^3 =(\frac{n(n+1)}{2})^2"